OpenStudy (anonymous):

A minute hand of a clock is 22cm long. Find the distance traveled by the tip of the minute hand in one hour. Please help, this is due tomorrow and could you please include the working out and methods to help me tackle these strategies in future.

OpenStudy (mrnood):

What is the formula for the circumference of a circle? (Hint - it has pi and r in it...)

OpenStudy (anonymous):

3.14 oor 22 over 7 obviously but still I don't get this.

OpenStudy (mrnood):

That is NOT the formula for the circumference of a circle - that is the value of pi. If oyu don't already know it then type "circumference of circle" into google and come back with the formula.

OpenStudy (anonymous):

oh great sorry i thought you meant definition

OpenStudy (anonymous):

i know that it is pi x d or pi x r x 2

OpenStudy (anonymous):

im called a math dunce for a reason i fear my math teachers

OpenStudy (mrnood):

OK - you have the correct formula. NOW - the question tells you that the hand is 22cm long - this means that its tip moves round a circle with radius 22cm. It moves once round the clock face in 1 hour so it travels a distance equal to the circumference of a circle 22cm radius. So you know the formula, and you know the radius - so you can work out the distance travelled (= circumference)

OpenStudy (anonymous):

right so is it 22 x 2 x pi? ok then if it is then it would be 138.16cm?

OpenStudy (mrnood):

I make it 138.23 but you are probably making an approximation ofr pi - so that's ok. You should fear your own lack of confidence - not the teachers! well done :-)

OpenStudy (anonymous):

but what if it is 40 or 30 mins how can i do it? going to high school next year and i'm a total dunce at almost anything math related

OpenStudy (mrnood):

well there are 60 mins in 1 hour so 40 mins = 40/60 of 1 hour = 2/3 hour. So the hand moves 2/3 * distance in general no of mins/ 60 * circumference

OpenStudy (mrnood):

OK - soz got to go, Good luck

OpenStudy (anonymous):

bye